OpenStudy (anonymous):

Question 7 (Multiple Choice Worth 1 points) [1.06] MA.8.A.6.1 Convert 10 to the twelveth power from exponential to standard form. 0.000000000001 0.0000000000001 100,000,000,000 1,000,000,000,000

OpenStudy (anonymous):

@amistre64 @Loser66 @Luigi0210 @reemii

OpenStudy (anonymous):

....

OpenStudy (reemii):

ah, "10 to the 12th power" is \(10^{12}\). that is , 100000000..0 (12 zeros).

OpenStudy (anonymous):

so its the last one?

OpenStudy (reemii):

i think so

OpenStudy (anonymous):

Thank you! :D

OpenStudy (anonymous):

Question 8 (Multiple Choice Worth 1 points) [1.07] MA.8.A.6.4 Divide the quantity eighty-one point nine times ten to the seventeenth divided by the quantity nine times ten to the negative thirteenth. Express the answer in scientific notation. nine point one times ten to the negative thirtieth nine point one times ten to the negative fourth nine point one times ten to the fourth nine point one times ten to the thirtieth

OpenStudy (reemii):

\(\frac{81.9\times 10^{17}}{9\times 10^{-13}} \) ?

OpenStudy (anonymous):

Yes, that's how my problem looks.

OpenStudy (anonymous):

..................

OpenStudy (reemii):

\(81.9/9 = 9.1\) \(\frac{10^{17}}{10^{-13}} = 10^{17+13}\).

OpenStudy (anonymous):

so its the last one ?

OpenStudy (reemii):

yes.

OpenStudy (anonymous):

Thanks again.

OpenStudy (anonymous):

Question 11 (Multiple Choice Worth 1 points) [1.02] MA.8.A.6.4 Which is not an integer? -3 0 3.0 -3.8